Zinc oxide, ZnO(s), is formed by the reaction of zinc sulphide, ZnS(s), with oxygen. 2ZnS(s) + 3O2(g) ---> 2ZnO(s) + 2SO2 If 16.7 g of zinc sulphide reacts with 6.70 g of oxygen, what mass of zinc oxide is produced?
1
Expert's answer
2013-05-09T07:10:33-0400
General formula: moles=mass/molar mass
If we let: n = moles of pure substance,
m = mass of substance in grams,
M = molar mass of the pure substance in g mol-1
we can write the equation: n=m/M
The molar mass of atoms of an element is given by the atomic weight of the element. The molar mass of a compound is given by the sum of the standard atomic weights of the atoms.

n(ZnS) = 16.7g /( 97 g/mol)n(ZnS) =0.172 mol
n(O2) = 6.7g/(32g/mol)=0.2 mol
n(ZnS)=n(ZnO), because from the balanced chemical equation the mole ratio ZnS:ZnO is 2:2

n (ZnO) = 0.172 mol
General equation can be rearranged to give the following:
mass=moles*molar mass

m (ZnO) = 0.172 mol * 72 g/mol = 12.384 g

Answer:
mass of zinc oxide=12,384 g